OpenStudy (anonymous):

Answer only please... Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach? 0.8 ft 1.6 ft 9.8 ft 19.6 ft

OpenStudy (anonymous):

Hi are you in physics? we did this in physics

OpenStudy (anonymous):

I got 31.8

OpenStudy (anonymous):

checked work but still got 31.8

OpenStudy (anonymous):

im in algebra 1 :P

OpenStudy (anonymous):

do i need to show my work?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i used the 5 kinematic equations. do you understand?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

the answer i got is not one of the choices though

OpenStudy (anonymous):

i got 0.8 idk if thats right though

OpenStudy (anonymous):

alright i used the kinematic equations: Vf=vi+a(t) and deltax=vi(t)+1/2a(t)^2

OpenStudy (anonymous):

its 9.8 at maximum height velcoity will be zero hence max h= 25X25/(2X32)=9.8